Consider a parallel plate capacitor which is maintained at potential 200V .
The separation distance between the plates of capacitor and the area of plates are 1mm and 20cm^(2) respectively.Calculate displacement current in 1microsecond .
ANS: 3.5mA
1
Expert's answer
2020-09-23T11:43:18-0400

Between the plates of the capacitor, there exists a potential difference, V of 200v. The plates are 1mm apart, and the area of the plates is "20cm^{2}"

Now, V= 20v, distance, d = 1mm = "1\times10^{-3}m"

A = 20cm"^{2} = 20 \times10^{-4}m^{2}" , and time, t in micro-seconds "\mu s = 10^{-6} s"

A change in electric flux and electric field produces displacement current given by the formula

"I_{d} = \frac{\epsilon_{o} d \Phi_{B}}{dt}" where "\Phi_{B}" is electric flux, and "I_{d}" is the displacement current.

"I_{d} =\frac{ \epsilon_{o} EA}{t}" where the value of "\epsilon_{o} = 8.85\times10^{-12}"

Electric field, E is given by E = "\frac{V}{d}"

Thus "I= \frac{VI_{d}}{d} = \frac{\epsilon_{o} VA}{td} = \frac{8.85\times10^{-12}\times200\times20\times10^{-4}}{10^{-6}\times1\times10^{-3}}"

"I = 35400 \times 10^{-7} = 3.5 mA"